Calculate the quantity demanded when: (0) P= N500, (i) P= N1000, (ii) PeN1500 and

(iv) P=N2000.


1
Expert's answer
2021-06-15T12:44:59-0400
"Q_D=a-bp"

(0) P= N500,


"Q_D=a-500b"

(i) P= N1000,


"Q_D=a-1000b"

(ii) PeN1500


"Q_D=a-1500b"

and

(iv) P=N2000.


"Q_D=a-2000b"